OpenStudy (anonymous):

simplify z^2-4/2z^2+5z+2 x 2z^2+z/z-2

OpenStudy (luigi0210):

It's this, correct? \[\LARGE \frac{z^2-4}{2z^2+5z+2}*\frac{2x^2+z}{z-2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no 2z^2+z on the right

OpenStudy (luigi0210):

Oh, sorry. But start off by factoring.

OpenStudy (anonymous):

these are my options 1. 2z^2/5z 2. 2z^2+z/2z+1 3. z 4. 1/z 5. none

OpenStudy (anonymous):

i dont need to show work on them

OpenStudy (luigi0210):

Still, it's a good thing to do so.. it helps when they get lengthy and complicated..

OpenStudy (anonymous):

aww man im kind of in a rush lol

OpenStudy (luigi0210):

Well okay.. first write it out: \[\LARGE \frac{z^2-4}{2z^2+5z+2}*\frac{2z^2+z}{z-2}\] Now factor: \[\LARGE\frac{(z-2)(z+2)}{(2z+1)(z+2)}*\frac{2(z+1)}{z-2}\] Now just cross cancel and simplify.. pretty simple

OpenStudy (luigi0210):

On the top right, that should be z(2z+1) not 2(z+1)

OpenStudy (anonymous):

thanks man

OpenStudy (luigi0210):

Anytime!